OpenStudy (anonymous):

A. Find the derivative of f at x. That is, find f'(x). B. Find the slope of the tangent line to the graph of f at each of the two values of x given to the right of the function f(x)=x^2-8; x=-1, x=3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hey

OpenStudy (misty1212):

you know the derivative of \(x^2\)?

OpenStudy (misty1212):

if the answer is "no" that is fine, i will tell you

OpenStudy (anonymous):

is it 2x?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

and the derivative of a constant is zero, making the derivative of \(f(x)=x^2-8\) the function \[f'(x)=2x\] that is all

OpenStudy (anonymous):

but its two parts to the question is 2x the answer to them both

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

no \(2x\) is the derivative you now need the equation for two lines, one where \(x=-1\) and the other where \(x=-3\)

OpenStudy (anonymous):

f(x)=-1^2-8 and f(x)=-3^2-8

OpenStudy (misty1212):

if \(x=-1\) then \(y=(-1)^2-8=-7\) so the point is \((-1,-7)\) and the slope is \(2\times -1=-2\) use the point slope formula to get the equation of the line

OpenStudy (misty1212):

if \(x=3\) then \(y=3^2-8=1\) the point is \((3,1)\) and the slope is \(2\times 3=6\) points slope again

OpenStudy (anonymous):

y-(-7)=-2(x-(-1)) and y-(1)=6(x-(3))

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

you can clean them up a great deal, but yes

OpenStudy (anonymous):

y=-2x+5 y=6x-17

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

o Thank you